# 3DGSR: Implicit Surface Reconstruction with 3D Gaussian Splatting

### v1

@@ 5.2. Ablation Study
We conduct ablation studies on the NeRF Synthetic dataset in Table 3 to evaluate the effectiveness of each module in our method. Six different configurations are investigated to train our model.
@@
The volumetric constraints have a minimal impact on novel view synthesis (PSNR: 32.83-33.32) yet significantly influence the reconstruction results (Chamfer-L1: 3.63-1.51). Both 3DGS rasterization and ray-based volumetric rendering rely on image supervision. However, the depth and normal representations derived from these dual methods exhibit distinct characteristics. While 3DGS excels at capturing high-frequency signals, it often suffers from floaters. Conversely, ray-based volumetric rendering yields smoother and more stable outcomes, but may struggle to capture fine details.
These two methods can complement each other and help address their respective weaknesses. Redundant Gaussian points can be effectively eliminated, and simultaneously, the high-frequency information learned by 3DGS can be employed to enhance the quality of the SDF field. This is demonstrated in Table 1 and Figure 3 of the main paper.
Learnable $\beta$.
By making $\beta$ trainable results, the conversion from SDF to 3DGS opacity can be learned automatically, which results in a 0.20 and 0.44 improvement in rendering and reconstruction performance, respectively.
Normal Supervision.
If we use the pseudo normal to supervise the normal of Gaussian points, the model will get 0.23 and 0.05 improvement in Chamfer-$\mathcal{L}_{1}$ and PSNR, respectively.
Eikonal Loss.
Removing the eikonal loss (Eq. (10)) will relax the constraint of SDF field, which can damage the surface learning and lead to worse reconstruction results. Note that better rendering quality (PSNR 33.4) is obtained in this configuration, which indicates the relaxation of geometric constraints in specific ways can aid in improving rendering quality.
Surface Alignment.
Constraining the Gaussian points close to the zero-level surface will inject more high-frequency information into the SDF field leading to detailed surface extracted results.
@@
Compared with the marching cube with the SDF field, we also employ Poisson reconstruction to extract the surface from Gaussians with normals derived from the implicit SDF field, which exhibits a lower performance. This can be attributed to the noisy, sparse distribution of Gaussians that significantly impact the surface reconstruction quality. In contrast, the implicit SDF field provides a continuous representation, allowing for easy extraction of high-fidelity surfaces.
Method Chamfer-$\mathcal{L}_{1}$ F-score PSNR LPIPS SSIM
w/o Geo 3.63 80.20 32.83 0.0380 0.963
static $\beta$ 1.95 90.20 33.02 0.0385 0.966
w/o eq (9) 1.74 91.36 33.18 0.0383 0.967
w/o eq (10) 2.00 87.89 33.40 0.0350 0.968
w/o eq (11) 1.69 91.70 33.06 0.0388 0.966
Poisson 2.03 87.86 - - -
Ours (full) 1.51 93.50 33.23 0.0340 0.968
Table 3. Ablation study on the NeRF Synthetic dataset, where we evaluate the effect of each proposed component.
5.3. Experimental Results
Tables 1 and 2 present the quantitative results for synthetic and real-world datasets, respectively. On synthetic data, Table 1 demonstrates that our method surpasses other state-of-the-art methods in terms of the reconstruction quality, as measured by Chamfer-$\mathcal{L}_{1}$ distance. Furthermore, our method also excels in synthesizing high-quality novel view images, which are comparable to state-of-the-art novel view synthesis methods. This highlights our method’s superiority over surface reconstruction techniques and its competitive performance in novel view synthesis tasks. On real-captured DTU datasets, as shown in Table 2, our method also exhibits the best performance in terms of average surface reconstruction quality.
Besides, we also found that better reconstruction results have less contribution to the rendering quality compared to the vanilla 3D-GS. The slightly lower performance is because 3DGS is optimized solely for rendering images; while our geometry constraints reduce the freedom of the optimization of 3DGS for the rendering tasks, the best solution for rendering may not necessarily be for geometry, as evidenced by performance in Table 1 (Trade-off between rendering quality and surface reconstruction). But overall, our formulation achieves the best trade-off between rendering quality and reconstruction quality.
Additionally, we provide qualitative comparisons of the reconstruction results qualitatively and present them in Figure 5. The state-of-the-art NeRF-based 3D reconstruction method NeRF2Mesh (Tang et al., 2023b) fails to preserve fine details like fine Mic textures, Lego tracks, and stripes on the Hotdog disk. RelightableGaussian (Gao et al., 2023) incorporates shading equations to determine Gaussian appearance and trains 3D Gaussians with attributes of normal and BRDF. 3D surfaces can then be obtained by applying Screened Poisson Reconstruction (Kazhdan and Hoppe, 2013) on the 3D Gaussians assigned with normal. However, Gaussians are naturally discontinuous and unstructured, resulting in broken parts on the Mic and floaters on the Hotdog disk. Due to the absence of smooth surface modeling, its reconstructed results on Lego tracks, which contain intricate structures, are noisy and inaccurate. Combining the fitting power of Gaussian Splatting and the smooth surface modeling of the implicit SDF field, our method reconstructs surfaces with geometric details accurately and smoothly.
@@ 6. Conclusion
In this paper, we introduce 3DGSR, an efficient method for high-quality surface reconstruction using 3DGS. Our approach is based on two key components: 1) the integration of neural implicit SDF and its binding with 3D Gaussians through a differentiable SDF-to-opacity transformation function, and 2) the utilization of volumetric rendering and the SDF-and-Gaussian geometry consistency regularization for SDF optimization. The first component allows for the joint optimization of SDF and 3D Gaussians, with the optimization of Gaussians providing supervision signals for learning the SDF. The second component supplies additional supervision signals for refining the SDF in areas not occupied by Gaussians, using a consistency loss that aligns depth (normal) from SDF with that derived from Gaussians. Our extensive experiments showcase the effectiveness of 3DGSR in reconstructing high-quality surfaces outperforming those obtained from state-of-the-art reconstruction pipelines, without compromising the rendering capabilities and efficiency of 3D Gaussians. Our study does have certain limitations caused by the trade-off between rendering quality and surface smoothness. In order to render high-quality images, the smoothness of the surface may be compromised in some cases involving complex texture patterns.

### v2

@@ 5.3. Ablation Studies
We conduct ablation studies on the NeRF Synthetic dataset in Table 5 to evaluate the effectiveness of each module in our method. Four different configurations are investigated to train our model. We also show the qualitative results in Fig. 10.
@@
The volumetric constraints have a minimal impact on novel view synthesis (PSNR: 32.83-33.86) but significantly influence the reconstruction results (Chamfer-$\mathcal{L}_{1}$: 3.63-1.37). This is because the volumetric rendering for depth and normal supervision can provide dense supervisory signals for learning the continuous SDF. Without these constraints, numerous noisy structures would be generated in areas not occupied by Gaussians, negatively affecting the overall reconstruction quality. As shown in Fig. 10, if we directly use the tight coupling strategy without any geometry constraint, we can only get the coarse result.
@@
Compared with the marching cubes on the SDF field for surface extraction, we also employ Poisson reconstruction to extract surface from Gaussians with normals derived from the implicit SDF, which yields a much lower performance. This demonstrates the advantage of using learned implicit SDF to represent the surface that is continuous, more tolerant to noisy supervisions, and allows for easy extraction of high-fidelity surfaces. Notably, the surface is modeled as a hash grid and a single-layer MLP, which is lightweight and does not introduce many burdens for computation.
Coupling Strategy.
We experiment with the tight coupling and loose coupling strategies proposed by our approach. For the tight coupling strategy, the learning of the parameter $\beta$ impacts our performance. In contrast, the loose coupling strategy leads to improved results in both reconstruction and novel view synthesis, demonstrating its effectiveness in balancing the different aspects of the modeling process.
Figure 7. Sparse point cloud and the influence of projected depth supervision $D_{sfm}$ in 2DGS and 3DGSR.
Projected Sparse Depth $D_{sfm}$.
Method $D_{sfm}$ Chamfer-$\mathcal{L}_{1}$
2DGS (Huang et al., 2024) Yes 0.85
2DGS (Huang et al., 2024) No 0.80
Ours Yes 0.70
Ours No 0.81
Table 6. Ablation study of sparse depth supervision. Yes indicates the use of $D_{sfm}$ as a supervision signal, while No indicates that $D_{sfm}$ is not used for supervision.
To ensure a fair comparison, we conducted experiments to analyze the influence of the supervision by projected sparse depth $D_{sfm}$ on both 3DGSR and 2DGS (Huang et al., 2024), presenting the quantitative results in Table 6 and qualitative results in Fig. 7. As shown in Table 6, we observed that adding sparse depth cues negatively affects the performance of 2DGS (Huang et al., 2024), causing a worsening in Chamfer-$\mathcal{L}_{1}$ from 0.80 to 0.85. This performance drop may stem from the fact that 2DGS (Huang et al., 2024) relies on depth fusion for reconstruction, as opposed to our approach, which jointly optimizes and aligns the surface with the 3D Gaussians. In contrast, the addition of sparse depth supervision proves beneficial in our method, resulting in improved quantitative results (Chamfer-$\mathcal{L}_{1}$ from 0.81 to 0.70). These cues assist SDF optimization in regions where image supervision is sparse and camera poses are inaccurate, though they are less impactful in areas where sufficient observations are available. When trained using only image signals, our method can be comparable with 2DGS (Huang et al., 2024). Fig. 7 further demonstrates that 2DGS (Huang et al., 2024) shows only minor improvements with sparse depth supervision, whereas our method achieves more detailed reconstructions even in the absence of such supervision. In conclusion, these depth cues are valuable for improving performance in cases of inaccurate camera poses and ill-posed sparse view data, especially in real-world scenes. Our approach can effectively utilize these free cues to enhance reconstruction quality.
@@ 6. Conclusion
In this paper, we introduce 3DGSR, an efficient method for high-quality surface reconstruction using 3DGS. Our approach is based on two key components: 1) the integration of neural implicit SDF and its alignment with 3DGS, and 2) the utilization of volumetric rendering and the SDF-and-Gaussian geometry consistency regularization for SDF optimization. The first component allows for the joint optimization of SDF and 3D Gaussians, with the optimization of Gaussians providing supervision signals for learning the SDF. The second component supplies additional supervision signals for refining the SDF in areas not occupied by Gaussians, using a consistency loss that aligns depth (normal) from SDF with that derived from Gaussians. Our extensive experiments showcase the effectiveness of 3DGSR in reconstructing high-quality surfaces outperforming those obtained from state-of-the-art reconstruction pipelines, without compromising the rendering capabilities and efficiency of 3D Gaussians. We hope our approach could inspire more future work. Despite its promising results, our approach still has limited capability to handle unbounded scenes, due to the use of a hash grid and implicit SDF, and also has worse results on real datasets with noisy pose estimation results. Furthermore, it may encounter difficulties in reconstructing transparent objects. These limitations highlight potential areas for further improvement and development in our method.
7. Acknowledgement
This work has been supported by the Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422, 17212923), Theme-based Research (Grant No. T45-701/22-R) and RGC Matching Fund Scheme (RMGS). Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust.